OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 5i

OpenStudy (anonymous):

f(x) = x4 - 6x3 - 20x2 + 122x - 928 f(x) = x4 - 19x2 + 244x - 928 f(x) = x4 - 6x3 + 20x2 - 122x + 928 f(x) = x4 - 61x2 + 244x - 928

OpenStudy (anonymous):

medals will be given btw plz help

OpenStudy (anonymous):

because of the given roots, multiply out (x-4) (x+8 )(x-[2+5i]) (x-[2-5i]) to get the polynomial you need

OpenStudy (anonymous):

so just use algebra to solve

OpenStudy (anonymous):

thanks @dpaInc

OpenStudy (anonymous):

ur welcome... sorry internet probs...